OpenStudy (anonymous):

Add: x+4/x^2-9 + 1/x-3

OpenStudy (anonymous):

\[\frac{x+4}{(x-3)(x+3)}+\frac{1}{x-3}\] common denominator is \((x+3)(x-3)\) \[=\frac{x+4}{(x-3)(x+3)}+\frac{x+3}{(x+3)(x-3)}\] now add straight across

OpenStudy (anonymous):

\[\frac{ x+7 }{ (x-3)(x+3) }\] ?

OpenStudy (anonymous):

@satellite73 ^